# Price Experimentation and Interference

### v4

@@ 9 Conclusion and Future Work
In this paper we studied the biases that arise due to interference in platform experiments, with a particular emphasis on profit and revenue metrics. Notably, we find that in certain settings, these biases can be so severe that they introduce a “change of sign” into the canonical estimators for profit, suggesting to firms that they should move prices (or fees) in the wrong direction. We then derive a novel method for debiasing canonical A/B experiments, by splitting units equally between treatment and control. We conclude by giving empirical evidence for the plausibility of estimators with the wrong sign in practical settings, arguing that firms should be more cautious when using canonical A/B pricing experiments.
In future work we aim to extend our debiasing result, to allow for the correction of canonical experiments in cases where the treatment/control split is not 50/50. In addition to being more flexible, this may allow us to mitigate the sampling inefficiency of our current debiasing method. Currently, our method only uses half of the experimental data, namely, only the observed control demand, which may be a source of increased variance. Finally, we aim to run empirical studies to get a better sense of the true prevalence of changes of sign in practice.

### v5

@@ 9 Conclusion and Future Work
In this paper we studied the biases that arise due to interference in platform pricing experiments, with a particular emphasis on profit and revenue metrics. Notably, we find that in certain settings, these biases can be so severe that they introduce a “change of sign” into the canonical estimators for profit, suggesting to firms that they should move prices (or fees) in the wrong direction. We then derive a novel method for debiasing canonical A/B experiments, by splitting units equally between treatment and control. We conclude by giving empirical evidence for the plausibility of estimators with the wrong sign in practical settings, arguing that firms should be more cautious when using canonical A/B pricing experiments.
We propose multiple direction for future work. First, we believe that more analysis needs to be done on the debiasing of canonical estimators. Our debiasing method, while novel, relies on the treatment/control split being precisely 50/50. A debiasing method that functions in all treatment split settings would be both more flexible, and would allow us to mitigate the sampling inefficiency of our current debiasing method. Our current debiasing approach only uses half of the available data, namely only the control demand, which may be a source of increased variance in practice.
A second, and related direction for future work involves quantifying the variance of these estimators themselves. As we note in Section 7, an interval of prices always exists in which a sign flip will occur in expectation (for a logit demand function), but this interval is bordered by the optimal price. In close-to-optimal settings, it is possible for estimators to be assigned the wrong sign entirely by fault of natural variance in the experiment. Typically, practitioners avoid this by dismissing results that are not statistically significant, meaning many of these incorrectly signed estimators may be ignored in practice because they fail to reach some critical p-value. Identifying conditions under which we would expect to see a statistically significant estimator with the wrong sign would be helpful to fully quantify how these sign flips actually impact practitioners. That said, our view is that practitioners should be more concerned with the impact of bias as studied in our paper, since variance vanishes with large sample sizes while bias does not.
Finally, our paper motivates the need for further empirical study into this phenomenon. Our empirical results suggest that estimators with the wrong sign may be more prevalent than previously expected. One approach might be for practitioners to employ a ”meta-experiment” similar to Holtz et al., (2025), with both a canonical LR experiment and a cluster-randomized experiment, as a way to compare the usual approach to a relatively less biased alternative. More generally, our work suggests there is benefit in developing empirical tests to identify when debiasing is needed to avoid making wrong decisions.